OpenStudy (bakonloverk):

-18x-20=-4y how do i find the slope and graph it?

OpenStudy (amistre64):

to graph it, you can find 2 points that satisfy the equation ... some obvious ones tend to be the x and y intercepts

OpenStudy (amistre64):

the slope can be defined as the y intercept divided by the x intercept

OpenStudy (bakonloverk):

that would make it 0.22

OpenStudy (amistre64):

perhaps ... -18x-20=-4y ; (0,5) -18x-20=-4y ; (-10/9,0) -5/(10/9) = -45/10 = -4.5

OpenStudy (amistre64):

missed something in there, im sure of it

OpenStudy (bakonloverk):

how did u get the (o,5) and (-10/9,0)?

OpenStudy (amistre64):

well, the y intercept is the value of y at x=0 the x intercept is the value of x at y=0

OpenStudy (amistre64):

-18(0) - 20 = -4y ... (0,5) -18x - 20 = -4(0) ... (-20/18, 0) , -10/9

OpenStudy (bakonloverk):

oh ok tht makes sense now

OpenStudy (amistre64):

\[slope=\frac{y~int}{-(x~int)}\]

OpenStudy (bakonloverk):

wait wait wait wait, is tht how i solve it?

OpenStudy (amistre64):

that is one way to solve it yes

OpenStudy (bakonloverk):

so is it -10/9?

OpenStudy (amistre64):

that is the x intercept

OpenStudy (bakonloverk):

ok this is confusing me, so what is the y?

OpenStudy (amistre64):

i solved 2 points ...

OpenStudy (amistre64):

0,5 and -10/9, 0 since the x is -10/9, the leaves .... 5 for the y

OpenStudy (bakonloverk):

oh ok i get it now, thank you

OpenStudy (amistre64):



OpenStudy (anonymous):

The way I would look at it without calculus, is you need to have a single y on the right side. Before that however, we want to find the slope. The slope can be found by looking at the coefficient of x after y is single (coefficient of 1) You may need to perform a few operations to get there (in this case, dividing both sides by 4 get just "y" on the right). After division: y=4.5x+5 We can see that the slope is m=4.5. Now to graph it, we want to find the y intercept. The y intercept is the value of y when x=0. This will let us know where the line crosses the y axis. when x=0, y=4.5(0)+5=5 The y intercept is -5, so keep that in the back of your head for now. Now we want to do the same for the x intercept, by setting y=0. when y=0, 0=4.5x+5 x=-10/9 To draw this, the line must intersect both the y axis at -5, and the x axis at -10/9. Since this isn't a quadratic equation (or anything more complex) it will be a straight line connecting those two points [(0,5) and (10/9,0)]. Sorry if I was late, and hope that helps!

OpenStudy (bakonloverk):

thank u soooo much

OpenStudy (anonymous):

Oops, sorry for the double typo, replace anywhere I wrote "-5" as the y intercept with +5!